### v2

# ElectricityEmissions.jl: A Framework for the Comparison of Carbon Intensity Signals

Abstract.
An increasing number of individuals, companies and organizations are interested in computing and minimizing the carbon emissions associated with their real-time electricity consumption. To achieve this, they require a carbon signal, i.e. a metric that defines the real-time carbon intensity of their electricity supply. Unfortunately, in a grid with multiple generation sources and multiple consumers, there is no unambiguous way to trace electricity from source to sink. This makes it hard to define an appropriate signal, leading to a raging discussion about how to best quantify the carbon footprint of electricity. This paper seeks to inform the discussion about which carbon signal is better or more suitable for two important use cases, namely carbon-informed load shifting and carbon accounting. We do this by developing a new software package ElectricityEmissions.jl, that computes several established and newly proposed carbon emission metrics for standard electric grid test cases. We also demonstrate how the package can be used to investigate the effects of using these metrics to guide load shifting. Our results affirm previous research, which showed that the choice of carbon emission metric has significant impact on shifting results and associated carbon emission reductions. In addition, we demonstrate the impact of load shifting on both the consumers that perform the shifting and consumers that do not. Disconcertingly, we observe that shifting according to common metrics such as average carbon emissions can reduce the amount of emissions allocated to data center, but cause an increase in the total emissions of the system.
1. Introduction
As the impacts on climate change become increasingly visible, many electricity consumers, from private persons to large corporations, show increasing interest in assessing and reducing their carbon footprint. For many consumers, a significant portion of their carbon footprint is tied to their electricity consumption and associated emissions from electricity generation. The source of electricity generation (e.g. solar, wind, hydro, nuclear, natural gas or coal) as well as the temporal availability of renewable power (i.e. daily and seasonal cycles of solar, wind and hydropower) vary widely between locations, leading to differences in carbon intensity of electric power across time and space. This provides an opportunity for consumers to reduce carbon emissions by using power where and when low-carbon electricity is available. However, to actively reduce the carbon emissions of their electricity supply, consumers need real-time information about how carbon emissions of electricity vary. Furthermore, they need a framework for leveraging this information not only for real-time load shifting but also for long-term carbon accounting, such that they can get credit for their efforts to reduce emissions. This is particularly important for companies and industries that are subject to emerging carbon emission regulations and reporting requirements, including e.g. producers of clean hydrogen (Internal Revenue Service, 2023).
Quantifying the carbon intensity of electricity consumption at differing times/grid locations, however, is not straightforward. While it is possible obtain information on emissions from the generators and assess how the total carbon footprint varies over time, the physics of the grid do not lend themselves to a clear definition of generation mix (and hence emissions intensity) at individual locations. This leaves significant confusion and open questions regarding what the best metric for carbon emissions, which is the topic of ongoing debate (Sukprasert et al., 2024). This paper develops a software package and an analysis framework to help evaluate the effectiveness of different carbon emission metrics.
1.1. Related work
Past research has focused on computing and analyzing different carbon emission metrics for real-time electricity consumption (e.g. (Electricitymaps, [n. d.]; Callaway et al., 2018; WattTime, [n. d.]; Lindberg et al., 2021b; Chen et al., 2023b, c)) or quantified impacts of current standards for long-term carbon accounting on long-term emissions (see (Langer et al., 2023) and references therein). Recently, there has also been increased interest in comparing these metrics for the purposes of carbon accounting and informing load-shifting applications. (Sukprasert et al., 2024; Richardson, 2023; Corradi, 2023)
The two main shifting metrics that are currently in use and being debated are ‘Average Carbon Emissions’ (ACE), defined as total system emissions divided by total system load, and ‘Locational Marginal Carbon Emissions’ (LMCE), defined as the emissions factor of the marginal generator(s) (i.e. the generator(s) that would respond if a load changes its consumption by a small amount). In a recent paper, the authors of (Sukprasert et al., 2024) compare marginal and average carbon emissions signals for the purpose of data-center workload scheduling. Their results, based on data from 65 regions throughout the world, found the two signals to be largely negatively-correlated, leading to significant differences in workload scheduling outcomes as well as accounted emissions.
Inter-metric variability in shifting-incentive and emissions accounting properties, such as that demonstrated in (Sukprasert et al., 2024), highlight the need for new tools and analysis frameworks to compare different emissions metrics. This is particularly important given that ACE and LMCE both have their drawbacks, which has led to interest in defining new metrics such as those based on ‘carbon flow’ methods (Chen et al., 2023a, b).
To inform the debate about the effectiveness and appropriateness of different carbon emission metrics, we develop a software package for comparing the different metrics from a power system perspective. Our comparison framework is based on emulating the market clearing mechanism of an electric grid (represented by solving a DC optimal power flow (OPF)) and computing total carbon emissions for the overall grid as well as various carbon emission metrics for different locations based on the result. This information can then be used for both carbon accounting (i.e. to allocate carbon emissions to different loads) and as an input to simulate real-time load-shifting, where a carbon-sensitive load, such as a data center, changes their consumption in response to the carbon metrics. Finally, it is possible to assess the effectiveness of real-time load shifting by re-running the simulation of the electricity market clearing with the updated electricity consumption pattern.
One important benefit of the framework is that it allows us (and other researchers) to evaluate both whether a given metric is effective in guiding real-time load shifting (i.e. if a consumer reduces their use of electricity in an hour with a high emission value and increases it in an hour with low emission value, does this lead to a reduction in overall system emissions?) and whether it possesses desirable properties for carbon accounting (e.g. does the metric guarantee that the total carbon emissions assigned to consumers are equal to the total carbon emissions from generation?). This can help inform the debate around current metrics. However, our second goal is to provide a platform for evaluation of newly proposed metrics. We demonstrate this by proposing a new carbon emission metric, namely the adjusted locational marginal emissions (ALMCE) metric, and assessing its merits relative to existing metrics using our framework.
1.2. Contributions
  • We provide a qualitative overview of existing carbon emission metrics and their pros and cons, and propose the new ALMCE metric to try to combine benefits of several metrics.
  • We contribute a new open source Julia package, ”ElectricityEmissions.jl” that enables easy calculation of several existing and emerging carbon emissions metrics, including the newly proposed ALMCE metric.
  • We demonstrate how the package can support quantitative comparison of different carbon emission metrics, both with and without load shifting, in a case study on the standard RTS-GMLC test system.
  • This case study is, to the best of our knowledge, the first to assess the impact of load shifting not only on total carbon emissions, but also on the accounted carbon emissions that are allocated to different types of loads.
2. Background
To explain the foundations of the real-time carbon accounting for electricity consumption, we start with a brief overview of the prevalent electricity market design in the United States, the emissions from generators and how changes in electricity consumption can impact these emissions.
Electricity Market Clearing
We consider an idealized setting, where the market clearing can be modelled mathematically as an optimal power flow (OPF) problem. The OPF problem minimizes the overall cost of operating the electric grid, subject to constraints that model the physical flows of electric power and limits on generation and transmission capacity. It can be expressed as a linear optimization problem. We provide the mathematical formulation for this optimization problem in Appendix A.
The OPF problem is solved approximately every 5-15 min (i.e. every time step $t$) in real-time operations of the electric grid. The main outcome of the OPF is the optimal generation dispatch, given by a set of generation setpoints $p_{G,t}^{*}$. The reason for repeatedly solving the OPF is to update the optimal dispatch $p_{G,t}^{*}$ in response to changes in the power system load $p_{D,t}$ and other system parameters, such as the time-varying availability of renewable generation represented by the time-varying maximum generation limits $p_{G,t}^{max}$.
Emissions from generators
Based on the optimal generation dispatch $p_{G,t}^{*}$, we compute the carbon emissions as
(1) $E^{tot}_{G,t}=\sum_{i\in\mathcal{G}}E_{G,i,t}(p_{G,i,t}^{*})=e_{G,i,t}p_{G,i,t}^{*}$
where $E_{G,i}(p_{G,i,t}^{*})$ is a function describing the emissions associated with generating the power $p_{G,i,t}^{*}$ from generator $i$. This function could take many forms (for example, generators may have different efficiency levels depending on how much they produce or the ambient temperature), but for practical reasons and lack of more detailed data, we assume that the carbon emissions per MW of generated power is constant, such that $E_{i}(p_{G,i}^{*})=e_{G,i}p_{G,i,t}^{*}$ where $e_{G,i}$ is the (constant) emissions factor of generator $i$.
Impact of changes in electricity consumption
In the OPF problem, increasing or decreasing consumption $p_{D,t+1}$ will change the optimal generation set-points $p_{G,t+1}^{*}$ and the resulting carbon footprint of the grid. This allows consumers to impact carbon emissions by changing their consumption, for example by shifting their consumption to a different location (e.g. by shifting computing loads within networks of data centers) or to a different time step (e.g. by deferring consumption until later in the day).
3. Carbon Emission Metrics
To define the carbon emissions of electricity consumption, we need to define a time- and location-specific carbon emission metric $e_{D,j,t}$ at each location in the grid, such that the carbon emissions of load $j$ at time $t$ given by
(2) $E_{D,j,t}=e_{D,j,t}p_{D,j,t}$
The challenge lies in choosing how to define $e_{D,j,t}$ (often also referred to as carbon intensity or emissions factor). Even though the carbon emission factor from each electric generator $i$ and time step $t$, denoted by $e_{G,i,t}$, are well defined, there is no clear consensus about how to allocate these carbon emissions to different consumers. This is in part because there is no unambiguous method to trace the actual flow of electricity from generator to load. Thus, multiple definitions of carbon emission metrics have been proposed. We summarize the definitions of some widely used metrics below, along with more recently proposed one and a brand new metric that we consider here for the first time. Note that this is not intended to be an exhaustive list of carbon metrics. Some other metrics, such as long-run marginal emissions (Hawkes, 2014), are not included below as they are less relevant to real-time carbon accounting and load shifting.
Average carbon emissions (ACE)
The most intuitive metric is the average carbon emissions $e_{D,t}^{avg}$, defined as the total carbon emissions $E_{tot,t}$ and divided by the total power consumption,
(3) $e_{D,j,t}^{avg}=e_{D,t}^{avg}=\textstyle{\frac{E^{tot}_{G,t}}{\sum_{j\in\mathcal{D}}p_{D,j,t}}}.$
With this definition, the emissions factor $e_{D,j,t}$ is the same for all loads $j$. As a metric, average carbon emissions has several benefits. It is easily accessible (i.e. published by system operators or estimated from publicly available information), is available for large geographical regions from data providers such as ElectricityMaps (Electricitymaps, [n. d.]) and WattTime (WattTime, [n. d.]), and can be used for scope 2 electricity accounting under the Greenhouse Gas Protocol (GHG Protocol, 2015). Average carbon emissions capture the emissions from all generators, and ensures that the total emissions from generation equal the total emissions assigned to loads. However, the average carbon emissions metric also has several drawbacks. It does not account for grid constraints that limit the transfer of power from one region to another. Further, average carbon emissions fails to capture how load shifting, i.e. changes in the timing or location of electricity consumption $p_{D,t}$, impacts total grid emissions, and does not reflect the availability of low-carbon generation sources, and, in particular, the presence of curtailed renewable energy.
Figure 1. Power generation profiles (top), average carbon emission (middle) and renewable curtailment (bottom) for June 10, 2021 in CAISO. We highlight hours with high curtailment (green) and low average carbon emissions (blue).
We illustrate these drawbacks based on an example from the CAISO grid (in California) on June 10, 2021. Fig. 1 shows the total generation for each fuel type (top), average carbon emissions measured in gCO2eq/MWh (middle) and amount of renewable curtailment in the grid (bottom). We observe that the average emissions are lowest in the late afternoon, when the renewable energy generation is highest. However, there is significant renewable energy curtailment earlier in the day, which is why the average emissions are higher in this time period. If the average emissions metric is used to shift load, consumers may choose to wait until later in the day with using power. However, from a grid perspective, this is opposite of what is needed, as such a shift would likely increase renewable energy curtailment and lead to overall higher emissions.
Locational marginal carbon emissions (LMCE)
To answer questions such as “how do carbon emissions change if I choose to charge my electric vehicle now rather than later?”, we can instead consider the marginal carbon emissions. By applying sensitivity analysis methods to the OPF problem, we can describe how a small change in a given load $\delta p_{D,j,t}$ changes the optimal generation dispatch $\delta p_{G,t}^{*}$. This allows us derive the associated change in emissions $\delta E=e_{G,t}^{T}\delta p_{G,t}^{*}$ and obtain the so called locational marginal carbon emissions (LMCE)
(4) $e_{D,j,t}^{lmce}=\textstyle{\frac{\delta E}{\delta p_{D,j,t}}}.$
Our implementation of LMCE is based on the method described in (Lindberg et al., 2021b), but extended to account for piece-wise linear cost functions. The full mathematical model is described in Appendix B. Note that LMCE inherently account for transmission constraints, and this varies by location in the grid when transmission capacity is limited.
The locational marginal emissions $\lambda_{CO2}$ are arguably the metric that most accurately describes the impact of (small) load shifting actions (i.e. intentional changes in $p_{D,t}$) to reduce carbon emissions (Lindberg et al., 2021a), (Lindberg et al., 2022). However, because they represent sensitivities of the OPF solution, LMCE can vary widely between time-steps and and may not be an appropriate metric for larger load shifts.
Further, it is unclear how LMCE can be used for carbon accounting, as it does not capture any information about emissions from non-marginal generators. Here, it is important to note the distinction between LMCE and locational marginal prices (LMPs), which are used to clear the electricity market. Since electricity markets seek to minimize total generation cost, the locational marginal price represent the marginal generation cost of the most expensive generator (i.e. the last one to get dispatched to serve electric load at a given location). Other generators that are running have a lower cost. For emissions, there is no such relationship, i.e. the marginal generator may have higher or lower emissions than the other units that are dispatched. As a result, if we allocate carbon emissions to loads based on the locational marginal emissions, the total allocated carbon emissions will generally be higher or lower than the total emissions from generation. In many cases we find that the difference between the total emissions from generation is significantly different from the total emissions allocated to loads.
To illustrate the volatility of LMCE, Fig. 2 shows the average and locational marginal emissions for ComEd in Chicago on 8/13/23. We observe that the 5 minute, real-time locational marginal emissions have much higher volatility than the hourly average emissions. We also see that the LMCE is significantly higher than the average emissions, indicating that the marginal generator at this location is more polluting than the average generator.
Figure 2. Marginal (blue) and average (orange) carbon emissions for Chicago (data from (PJM, [n. d.])).
Carbon flow-based metrics
Another class of metrics that have been considered in the academic literature are metrics based on a “carbon flow” which “traces” carbon emissions from generators to load (Kang et al., 2015; Chen et al., 2023c, b). The key idea behind these metrics is that generators inject both power and carbon emissions into the grid, and the carbon “flows” along with the electricity. The methods in (Kang et al., 2015; Chen et al., 2023c, b) are based on the so-called proportional power sharing rule (Bialek, 1996; Kirschen and Strbac, 1999), an assumption which stipulates that the power flowsinto a bus (and the carbon emissions flowing with it) are shared proportionally across the power flows out of it. It is important to note that this an assumption, not a physical law. There exist a variety of other proposals for tracing power flow (see e.g. (Chen and Dhople, 2019) and references therein), which would likely lead to different carbon allocations across loads. To the knowledge of the authors, carbon flow metrics have not been adopted in practice, possibly because their definition is less intuitive and more mathematically involved, and because the computation requires detailed knowledge regarding the network topology and flow of power. Nevertheless, these methods are interesting as they give rise to a range of possible carbon emission metrics, including locational average carbon emissions (Chen et al., 2023b).
Adjusted Locational Marginal Carbon Emissions (ALMCE)
Marginal metrics such as LMCE are appealing for use in informing load-shifting decisions, as they provide information on how grid emissions will change given a small change in load. However, performing carbon accounting using LMCE is problematic due to the lack of a guarantee that LMCE-assigned emissions will sum to actual total system emissions. To address this issue, we propose a new metric, the adjusted locational marginal carbon emissions (ALMCE). The core idea of ALMCE is to augment LMCE with an adjustment factor, which serves to normalize the sum of assigned emissions to be equal to the actual system emissions total. Mathematically, ALMCE is defined for load $j$ at time $t$ as follows.
(5) $e_{D,j,t}^{ame}=e_{D,j,t}^{lme}+\frac{E_{G,t}^{tot}-E_{D,t}^{lme}}{\sum_{j\in\mathcal{D}}p_{D,j,t}},$
where the first term is the locational marginal emissions $e_{D,j,t}^{lme}$ and second term represents the adjustment. This adjustment is defined as the difference between the total emissions from generation $E_{G,t}^{tot}$ and the total emissions assigned to loads using the LMCE metric $E_{D,t}^{lmce}=\sum_{j\in\mathcal{D}}e_{D,j,t}^{lmce}p_{D,j,t}$, shared across all loads proportional to their consumption.
While the full properties of ALMCE for carbon accounting and load-shifting remain to be determined through further analysis empirical testing, there are some properties worth noting. First, if there is a single marginal generator in the system, ALMCE is equivalent to ACE and in this case would lead to similar shifting results as ACE. Second, if we consider a single time step and sort nodes according to their nodal carbon emission values (e.g. from high to low), then ALMCE and LMCE would lead to the same ordering of nodes. Therefore, if we only consider spatial load shifting (where you can choose between consuming at different locations rather than at different times), ALMCE and LMCE might lead to similar shifting results.
4. ElectricityEmissions.jl
We have created a new Julia package, ElectricityEmissions.jl  [^1] Link to package, removed for doubleblind review , to serve as a central implementation of the carbon metrics mentioned in the previous section. This package is open-source, and contributions of other carbon metrics and accounting methods are welcome. Below is an example of computing and plotting carbon metrics for a small five-bus network.
Code Block 1 Calculating and plotting carbon intensity metrics
using PowerModels, ElectricityEmissions
using HiGHS
# Load the network data with generator emissions intensity
network = PowerModels.parse_file("case5_gen_intensity.m")
# Calculate carbon intensities (solves a PWL DC OPF problem)
lmce = calculate_LMCE(network, HiGHS.Optimizer)
almce = calculate_ALMCE(network, HiGHS.Optimizer)
ace = calculate_ACE(network, HiGHS.Optimizer)
lace = calculate_LACE(network, HiGHS.Optimizer)
# Calculate total system emissions
total_emissions = calculate_system_emissions(network, HiGHS.Optimizer)
# Plot network with carbon intensity data
# Update network data with chosen emissions intensity
update_emissions_intensity!(network,lmce)
# Plot the network
plot_emissions(network)
4.1. Carbon Intensity Calculation
The first part of the code example demonstrates how to compute the carbon intensity metrics. The network data and generator emissions information is loaded using the parse_file function from the PowerModels.jl package. It is passed to the functions calculate_ACE, calculate_LMCE, calculate_LACE and calculate_ALMCE to calculate the respective carbon intensity metrics. Each of these functions solve a piece-wise linear DCOPF problem to dispatch the system. Then load demand, optimal generator set-points, and generatorcarbon intensity (and for some metrics, the network structure) are used to calculate the nodal carbon intensity for each node in the network, as described in Section 3. We calculate total system emissions with the function calculate_system_emissions, which only needs generator set-points as an input.
4.2. Plotting
Visualization of the carbon metrics is important to understand the geographic spread of emissions intensity. To help users analyze their results, the functions update_emissions_intensity!() and plot_emissions() provide convenient data-processing and default settings for use with the power grid plotting package PowerPlots.jl (Wisconsin Power Optimization Lab ([n. d.]), WISPO-POP). An example of a visualization of the metrics is shown in the case study in Fig. 3.
4.3. Example Workflow
Here we step through an example workflow of using ElectricityEmissions.jl to estimate emissions and shift load, with counter-factual analysis.
1. Initial Emissions Calculations Prior to any load-shifting, use ElectricityEmissions.jl to calculate the initial carbon intensity at each bus in the system using the relevant calculate_...() function for each of the carbon intensity metrics being compared. These intensity values will have units of Tons CO2/MWh. To calculate assigned emissions (in Tons CO2) based on a particular metric, simply multiply the load at each bus by the corresponding carbon intensity value. Finally, store network intensities and assigned emissions according to each metric, as well as total system emissions (using calculate_system_emissions).
2. Solve a Load-Shifting Problem Use the carbon intensity values generated in the previous step (or outside carbon metric values/forecasts) as input to solve a load shifting optimization problem. Do this for each intensity metric and record the resulting post-shift loads. To estimate the impact of this shifting on the accounted carbon emissions, multiplying the initial carbon intensities by the new load values. Note that these are just estimates, for two main reasons. First, the carbon intensities might change as a result of the load shifting, leading to differences between the estimated and realized accounted carbon emissions. Second, the carbon intensities are themselves estimates of the carbon footprint of electricity. Although a load shifting action leads to a significant reduction in the accounted carbon emissions, it is not guaranteed that the total system emissions would actually decrease by the same amount.
3. Post-Shift Emissions Calculation Shifting load will result in new generator set-points and a new power flow which may change the carbon intensity values at each bus. To determine the final post-shift emissions, we rerun the relevant calculate_...() functions with the shifted load profile to recalculate the carbon intensity metrics. As explained above, the carbon intensities may change as a result of the load shifting. Using the post-shift carbon intensities, it is possible to compute for the accounted emissions that would be realized if the load-shifting was implemented. We refer to those as the realized accounted emissions.
We can also calculate the total post-shift system emissions, and compare those with the initial system emissions to assess the effectiveness of the load shift in reducing total carbon emissions.
5. Case Study: Accounting
We first provide a case study on carbon accounting according to the different carbon emission metrics, leveraging the ElectricityEmissions.jl package for computations.
5.1. Dataset
We make use of the RTS-GMLC (Barrows et al., 2020) test case, which contains 73 buses, 120 lines, and 158 generators. One year’s worth of hourly load and generation data is available for this system, totaling 8,784 cases. We make two modifications/additions, similar to the changes made to the same system in (Lindberg et al., 2021a). First, we add data-centers to the system by setting buses 103, 107, 204, and 322 to be data-centers, each with a nominal load of 250MW. Second, we use values from the U.S. Department of Energy(Office of Energy Policy and Systems Analysis, U.S. Department of Energy, 2016) to assign the following emissions intensities (in metric tons of CO2 per MWh) to each generator in the system:
  • Wind, Solar, Hydro: 0
  • Natural Gas: 0.6042
  • Oil: 0.7434
  • Coal: 0.9606
The data and code are available as examples in the package documentation  [^2] https://github.com/WISPO-POP/ElectricityEmissions.jl.
5.2. Methodology
We perform carbon accounting based on the LMCE, ALMCE, ACE, and LACE carbon intensity metrics across all time-steps in the dataset. ElectricityEmissions.jl is used to calculate nodal carbon intensities for all metrics, as well as total system emissions. Following this, accounted emissions are calculated for each metric through multiplication by nodal loads.
5.3. Results
Figure 3. Annual average nodal carbon emissions intensity, by metric. The buses are shown as circles, and generators are small squares arrayed around the buses. The color of the generators indicate the carbon emissions intensity of the source, while the color of the buses show the annual average carbon emissions. Locations of data centers are circled in red. Transmission lines connecting buses are shown in grey.
Carbon Emissions Intensity
(Tons CO2/MWh)
Whole
System
DC 103 DC 107 DC 204 DC 322
Metric Mean SD Mean SD Mean SD Mean SD Mean SD
LMCE 0.740 0.276 0.768 0.241 0.711 0.413 0.873 0.346 0.695 0.262
ALMCE 0.338 0.251 0.366 0.185 0.309 0.424 0.471 0.356 0.293 0.258
ACE 0.342 0.170 0.342 0.170 0.342 0.170 0.342 0.170 0.342 0.170
LACE 0.264 0.305 0.263 0.184 0.387 0.198 0.525 0.228 0.058 0.067
Table 1. Whole-Year Emissions Intensity Data
Carbon Emissions
(Million Tons CO2/MWh)
True
System
System
Acct.
DC
Acct.
DC 103
Acct.
DC 107
Acct.
DC 204
Acct.
DC 322
Acct.
Metric Sum Sum Sum Sum Sum Sum Sum
LMCE 15.828 33.012 6.692 1.686 1.562 1.917 1.527
ALMCE 15.828 3.162 0.803 0.679 1.035 0.644
ACE 15.828 3.008 0.752 0.752 0.752 0.752
LACE 15.828 2.707 0.577 0.850 1.153 0.126
Table 2. Whole-Year Emissions Accounting Data
Accounted emissions with different metrics
We first consider how the choice of carbon intensity metric impacts the total accounted carbon emissions across the year. Table 2 contains the accounted emissions for the entire system and each data-center location, as well as the true system emissions as measured by the generator set-points. In looking at the emissions assigned to the entire system, we see that ALMCE, ACE, and LACE leads to accounted emissions that equal the true system emissions value of 15.8 million tons of CO2. LMCE, however, lead to much higher accounted emissions at 33.0 million tons. This trend of much higher emissions with LMCE also holds true for the different data center locations.
This discrepancy happens because the marginal generators in the RTS-GMLC are much more polluting than the average generation source, and demonstrates the challenges of doing carbon accounting with LMCE.
Carbon intensities with different metrics
We next examining how the various carbon metrics assign emissions intensities to different system loads. Table 1 shows the emissions intensity statistics across the whole system, and for each data-center load individually. We show both the mean value across all hours of the year, as well as the standard deviation of the emissions intensity. First, we analyze inter-metric intensity differences for the system as a whole. On average, LMCE has the highest intensity values across the whole system, with a mean value of 0.740 Tons CO2/MWh. This is dramatically higher than the next-highest, ACE, with a mean intensity of 0.342 Tons CO2/MWh. ACE shows the least whole-system variability out of all the metrics, likely due to the fact that it does not vary spatially within a given time-step.
Second, we analyze emissions intensities across the four data-centers, located at buses 103, 107, 204, and 322. As in the whole-system statistics, LMCE assigns the highest average intensity values of any metric to all data-centers. However, the relative magnitude of ALMCE, ACE, and LACE varies greatly between location. Out of these three metrics, LACE has the highest mean value for the data-centers at buses 107 and 204, but the lowest for those at buses 103 and 322. Despite this variability, all of the metrics (save ACE, which is location-agnostic) are in agreement on the ranking of average emissions intensity between data-center locations. DC 204 has the highest emissions intensity, followed by DC 103, DC 107, and finally DC 322.
Finally, we consider the variation in mean carbon intensity across the full power system. Figure 3 shows the mean nodal emissions intensity values for each metric, with the data-center locations circled in red. We again see that LMCE (left) assigns by far the highest emissions, and that ACE (right) has the same values for all nodes. In this system, LACE shows the highest geographical variability in mean intensity values.
6. Case Study: Shifting
We next use the ElectricityEmissions.jl package and the example workflow set out in Section 4 to analyze what happens when the data centers use different carbon intensity metrics to guide spatio-temporal load shifting (i.e. shifting both between locations and across time steps).
6.1. Dataset
We make use of the same modified RTS-GMLC test case as described in the accounting-only case study, which provides a year’s worth of hourly load and generation data. Whereas the accounting-only case study added fixed 250MW data-center loads, the data-centers are new allowed to vary their load in order to minimize emissions.
6.2. Methodology
We perform load shifting based on LMCE, ALMCE, ACE, and LACE carbon metrics, within each 24-hour period of the dataset. Following the workflow described in Section 4.3, we first calculate initial ”pre-shift” emissions intensity for each metric.
Next, the data-centers, which are assumed to operate in a network controlled by a single entity, are then allowed to shift their load both in space and time to align with periods/locations of low carbon intensity. The mathematical formulations of the data center load shifting problem can be found in Appendix C. As an outcome of the optimization problem, we obtain estimated accounted emissions for the data centers.
Finally, the load-shifting step is followed by a re-dispatch of the system and recalculation of carbon intensity metrics in order to determine realized accounted emissions and whole-system emissions outcomes.
6.3. Results
Carbon Emissions (Million Tons CO2)
Pre-Shift Accounted Emissions
Post-Shift Accounted Emissions
Post-Shift Accounted Emissions
Actual Post-Shift
Emissions
Metric
Whole
System
DC
Loads
Non-DC
Loads
DC
Loads
Whole
System
DC
Loads
Non-DC
Loads
Whole
System
LMCE 33.011 6.692 26.320 6.363 (-4.19%) 32.689 (-0.98%) 6.632 (-0.90%) 26.057 (-1.00%) 15.669 (-1.00%)
ALMCE 15.828 3.161 12.666 2.865 (-9.36%) 15.833 (+0.03%) 3.015 (-4.62%) 12.818 (+1.20%) 15.833 (+0.03%)
ACE 15.828 3.008 12.820 2.824 (-6.12%) 15.880 (+0.33%) 2.896 (-3.72%) 12.985 (+1.29%) 15.880 (+0.33%)
LACE 15.828 2.707 13.121 2.370 (-12.45%) 15.825 (-0.02%) 2.595 (-4.14%) 13.230 (+0.83%) 15.825 (-0.02%)
Table 3. Shifting Case Study: Emissions Results
Figure 4. Change in Emissions Resulting From Spatio-Temporal Shifting
We analyze the emissions outcomes of the load shifting case study to understanding the following for all choices of shifting metric:
  • The effect of shifting on accounted carbon, as well as on the underlying carbon metric (ie its stability)
  • The differential impact of shifting on carbon responsibility assigned to shifters vs non-shifters
  • The effect of shifting on ground truth system emissions
  • The effect of shifting with one metric and accounting with another
Whole-year results
The results of the data center (DC) load-shifting case study are summarized in Table 3, and are divided into several categories.
Starting on the left, we show the pre-shift accounted emissions which are the carbon accounting values prior to any shifting (i.e. they are the same values as in the accounting case study, summarized here for convenience of the reader). Post-shift accounted emissions are divided into ”estimated” and ”realized”.
Next we show the estimated post-shift accounted emissions, which refer to the data-centers’ belief about what their accounted emissions will be after shifting, based on their new (post-shift) load values and the old (pre-shift) carbon metric values that they were optimizing for. We only show estimated post-shift accounted emissions for the data centers themselves, as these are the only loads that perform load shifting and obtain such estimates.
The realized post-shift emissions are the actual accounted emissions values, determined via post-shift loading and post-shift (i.e. re-calculated) carbon intensity values.
Finally, we show the total post-shift system emissions, both as an absolute number and a percentage difference to the no-shift total system emissions.
Impact of shifting on accounted emissions for data center loads
We begin by investigating whether load-shifting leads to a reduction in the accounted emissions assigned to data-center loads. To analyze this, we first analyze the realized accounted emissions for DC loads. We observe that shifting according to any of the metrics leads to a decrease the accounted emissions for DC loads, but to varying degrees. The greatest reduction in accounted emissions was achieved by ALMCE, folloed by LACE and ACE. The smallest reduction in accounted emissions happened when the data centers shifted according to LMCE.
Next, comparing the realized accounted emissions for the estimated accounted emissions, we observe a significant discrepancies. Notably, the realized reduction is significantly smaller than the estimated reduction (prior to the final post-shift recalculation of carbon metrics values). Estimated emissions reductions for ALMCE and ACE were  2x higher than realized, and  4x and  3x higher for LMCE and LACE respectively. These discrepancies indicate that the metrics changed due to load shifting, and in ways that reduced the overall effectiveness of the load shifting.
Impact of shifting on total system emissions and accounted emissions for non-DC loads
Next, we discuss how load shifting with the different carbon intensity metrics changes the overall system emissions, as well as the carbon emissions assigned to non-DC loads. We first observe that the only metric that significantly reduces the actual post-shift emissions is LMCE. For this metric, we observe a reduction of approximately 1% in accounted emissions for both non-DC loads and the overall system (similar to the reduction for DC loads). This matches the 1% reduction in total system emissions, even though the accounted emissions with LMCE are approximately 2x higher than the actual emissions.
In contrast, shifting with ALMCE or LACE barely changes total system emissions, while shifting according to ACE actually increases emissions by 0.33%. These observations hold true for both the total accounted emissions and the total emissions, as the metrics are designed to ensure that they match. Since the total system emissions hold steady (or increase) for these metrics while fewer emissions are assigned to DC loads, the accounted emissions for non-DC loads increased by 1̃%.
In conclusion, these results indicate that even though data centers are able to claim emission reductions after shifting with ALMCE, ACE and LACE, these emission reduction come from reallocating emissions to other loads rather than from actually decreasing emissions. In contrast, when shifting with LMCE, the data centers actually reduced total system emissions.
Impact of shifting on individual days
While over the course of the whole year, LMCE/LACE-based shifting seemed to decrease emissions and ALMCE/ACE-based shifting seemed to increase them, there is a large amount of variability in the performance of each metric on any particular day. Figure 4 shows the distribution of the daily changes in system emissions resulting from shifting with the different metrics, as well as the mean, median, 10th percentile and 90th percentile of the reductions. We can clearly see that shifting based on LMCE is much more likely to lead to large carbon emission reductions (i.e. large negative values) and less likely to increase the total carbon emissions from the overall system (i.e. has fewer positive values). We can also clearly see how some of the other metrics have a tendency to increase carbon emissions rather than reducing them.
Figure 5. Cross-Metric Shifting/Accounting: Change in Accounted Data-Center Emissions
Impact of shifting with one metric and accounting with another
Finally, we briefly examine the outcome of shifting based on one metric and accounting using another (e.g. determining data-center loads with LMCE, then assigning emissions using ACE). Figure 5 shows the absolute change in assigned data-center emissions, defined as the difference between pre- and post-shift accounted emissions (measured by the accounting metric).
As is to be expected, when the same metric is used for shfiting and accounting, we see the greatest reduction in accounted emissions. There is great variation in the ’cross-metric’ behavior of different ”shifting metric”-”accounting metric” pairs. For many of them, we see that shifting with one metric (e.g. LMCE) increases the accounted carbon emissions when a different metric is used for accounting. This indicates that companies or organizations who want to not only reduce overall carbon emissions, but also the carbon emissions that are allocated to them, have strong incentives to shift according to the same metric that is used for accounting.
7. Conclusions
Currently, there exists multiple competing definitions of the carbon intensity of electricity. In this paper, we seek to analyze benefits and drawbacks of existing, emerging and newly proposed carbon intensity metrics for two important use cases, namely load shifting and carbon accounting. To achieve this goal, we provide an overview and qualitative comparison of different metrics, and implement them in a new software package, ElectricityEmissions.jl. This software package provides tools for the calculation and analysis of several carbon emissions intensity metrics.
Leveraging this package, we compare the two most established carbon intensity metrics, namely average carbon emissions (ACE) and locational marginal carbon emissions (LMCE), against the recently emerging locational average carbon emissions (LACE) and a new metric proposed in this paper, which we refer to as adjusted locational marginal carbon emissions (ALMCE). We analyze the outcomes of using those metrics both in the context of accounting and load shifting.
In our case study, we find that load shifting based on the ACE, ALMCE and LACE intensity metrics has very little impact on overall carbon emission and instead tends to shift emissions from the data centers (which actively seek to minimize their emissions according to these metrics) to non-data center loads. In contrast, shifting according to LMCE reduces emissions for both data centers, non-data centers and the overall system. However, it is challenging to do accounting with LMCE, as the allocated carbon emissions are more than two times higher than the actual carbon emissions of the system.
The above findings are specific to the case we study, and further work is required to assess the effectiveness of the different metrics across a larger range of case studies. However, our results demonstrate that load shifting is not always a productive method to reduce short-term carbon emissions, and can be counter-productive if the carbon intensity metric is not chosen carefully.
In future work, we plan to expand the package to include other metrics and other functionality, allowing input from the community. We also wish to design new metrics with nice properties for both accounting and load shifting, and perform more thorough analysis of the differences between metrics on more realistic power system test cases and more realistic load-shifting setups.
Appendix A DCOPF Formulation
We use the following DCOPF formulation with piecewise linear costs to dispatch generation in ElectricityEmissions.jl, and for the linear generation sensitivity calculations necessary for the LMCE and ALMCE carbon metrics:
(6a) $\min_{\theta,p_{G},C_{G}}$ $\sum_{j}{C_{G,j}}$
(6b) s.t. $C_{G,j}\geq f_{j,k}(p_{G,j})$ $\forall j=1...N_{G},k=1...N_{ct}$
(6c) $\sum_{l\in G_{i}}p_{G,l}-\!\!\!\sum_{l\in D_{i}}p_{D,l}=\!\!\!\!\!\!\sum_{j:(i,j)\in L}\!\!\!\!\!\!-\beta_{ij}(\theta_{i}\!-\!\theta_{j})$ $\forall i=1...N$
(6d) $-P_{ij}^{lim}\leq-\beta_{ij}(\theta_{i}-\theta_{j})\leq P_{ij}^{lim}$ $\forall(i,j)\in L$
(6e) $P_{G,j}^{min}\leq p_{G,j}\leq P_{G,j}^{max}$ $\forall j=1...N_{g}$
(6f) $\theta_{ref}=0$
Here, the decision variables $\theta$, $p_{G}$, and $C_{G}$ refer to nodal voltage angles, generator set-points, and generator costs respectively. There are $N$ buses in the system and $N_{G}$ generators. The system is connected by lines $(i,j)\in L$, specifying the start and end buses $i$ and $j$.
Each generator has a $N_{ct}$-term piecewise-linear cost function. Equation 6b constrains the cost of each generator $j$ to be above all of the $k$ linear functions ($f_{j,k}$) that define its segments. Equation 6c is the standard nodal power conservation constraint, referencing generation ($G_{i}$), loads ($D_{i})$, and connected lines $(i,j)\in L$ at bus $i$. Equation 6d symmetrically limits the flow on each line $(i,j)$ by $P_{i,j}^{lim}$, and Equation 6e sets limits for each generator $j$ to be between $P_{G,j}^{min}$ and $P_{G,j}^{max}$. Finally, Equation 6f sets the voltage angle of the reference bus to zero.
Appendix B LMCE Calculation
LMCE is calculated as a linear sensitivity around a DCOPF solution, determining the effect on generation (and therefore emissions) of a single-unit change in load. For ElectricityEmissions.jl, we use the general method set forth in (Lindberg et al., 2021a) to calculate this sensitivity. However, we make several small modifications in order to deal with new constraints introduced by the piecewise-linear cost functions used in our DCOPF formulation.
The DCOPF formulation, as described in Appendix A, has $n=N+2N_{G}$ decision variables, where $N$ is the number of buses and $N_{G}$ is the number of generators. For each bus, the voltage angle $\Theta_{i}$ must be determined, and for each generator $j$, the active power generation $p_{G,j}$ and cost $C_{G,j}$ must be determined.
In the non-degenerate case, this means that the following linear system can be formed
(7) $Ax^{*}=b$
where $A\in\mathbb{R}^{n\times n}$ contains the problem’s equality and binding inequality constraints, $x^{*}\in\mathbb{R}^{n}$ contains the decision variable values at optimality, and $b\in\mathbb{R}^{n}$ contains the standard-form RHS constraint bounds.
As we are interested in the relationship between changes in load and changes in generation dispatch, we instead make use of the system
(8) $A\Delta x=\Delta b$
which is implied by Eq. 7. The constraints and variables are input to the system such that
(9) $A\begin{bmatrix}\Delta\Theta\\ \Delta P_{G}\\ \Delta C_{G}\end{bmatrix}=\begin{bmatrix}\Delta P_{D}\\ 0\end{bmatrix}$
where $\Theta\in\mathbb{R}^{N}$ refers to the vector of bus voltage angles, $P_{G}\in\mathbb{R}^{N_{G}}$ the vector of generator set-points, and $C_{g}\in\mathbb{R}^{N_{G}}$ the generator cost values. $P_{D}\in\mathbb{R}^{N}$ is the vector containing total load at each bus.
In this formulation, the nodal power balance constraints make up the first $N$ rows of the $A$ matrix (the only constraints to include $P_{D}$ variables). For ease in calculation, they are ordered according to their bus index. The final $n-N$ rows of $A$ correspond to the slack bus equality constraints, and all binding inequality constraints (for which the order does not matter).
We then invert the $A$ matrix, and solve for $\Delta P_{D}$. For compactness, we define the $B$ matrix to be the first $N$ columns and rows $N+1$ through $N+N_{g}$ of $A^{-1}$ These are simply the columns of $A^{-1}$ aligning with $\Delta P_{D}$ and the rows aligning with $\Delta P_{G}$ We can then write:
(10) $\Delta P_{G}=B\Delta P_{D}$
To determine the change in generation resulting from one unit of additional load at bus $i$ ($\Delta p_{G,i}$), we can set $\Delta P_{D}$ to be the ”one-hot” encoding of the bus in question (setting the $i$th entry to 1 and all others to 0). We can then perform the final LMCE calculation:
(11) $LMCE_{i}=\langle\Delta p_{G,i},e_{g}\rangle$
[…]
Appendix C Load Shifting Formulation
We perform spatio-temporal shifting, in which a group of data-centers are allowed to shift their load both to other data-center locations and other times of day. The spatio-temporal shifting problem is formulated as follows:
(12) $min_{d_{i,t}}$ $\sum_{i,t}{e_{D,i,t}p_{D,i,t}}$
(13) $s.t.$ $\sum_{i,t}{p_{D,i,t}}=N_{d}\cdot N_{t}\cdot D_{nom}$
$(1-\epsilon)\cdot D_{nom}\leq d_{i,t}\leq(1+\epsilon)\cdot D_{nom},$
(14) $\qquad\qquad\qquad\qquad\qquad\forall i=1....N_{d},t=1....N_{t},$
Here, $p_{D,i,t}$ and $e_{D,i,t}$ represent the load and carbon intensity at data-center $i$ and hour $t$ of the day. $N_{d}$ and $N_{t}$ are the number of data-centers and number of time-steps, in this case set to 4 and 24 respectively. $D_{nom}$ is the nominal data-center load (here set to 250MW), and $\epsilon$ is the data-center load flexibility (here set to 0.2). Eq. (13) enforces that the total data-center energy consumption (across all locations and time-steps) is the same as if all data-centers were at nominal load for the entire period. Eq. (14) defines a feasible range within which data center load can be adjusted up or down at a particular time-step (where the size of this range is controlled by $\epsilon$) . For our choice of parameters (4 data-centers with a constant nominal load of 250MW over 24 hours), the data centers have a total daily energy consumption of 24, [^0] MWh.
Acknowledgements.
This work is funded through the National Science Foundation, under awards #2328160 and #2325956.